Exercise 11-3 Stefani Company has gathered the following information about its product Direct materials: Each unit of product contains 3.20 pounds of materials. The average waste and spoilage per unit produced under normal conditions is 0.10 pounds. Materials cost $2 per pound, but Stefani always takes the 5.00% cash discount all of its suppliers offer. Freight costs average $0.25 per pound. Direct labor. Each unit requires 2.10 hours of labor. Setup, cleanup, and downtime average 0.10 hours per unit. The average hourly pay rate of Stefanis employees is $13.20. Payroll taxes and fringe benefits are an additional $2.20 per hour Manufacturing overhead. Overhead is applied at a rate of $4.80 per direct labor hour. Compute Stefanis total standard cost per unit. (Round answer to 2 decimal places, e.g. 1.25.) Total standard cost per units Click if you would like to Show Work for this question: Open Show Work LINK TO TEXT VIDEO: SIMILAR EXERCISE Question Attempts: 0 of 5 used SAVE FOR LATER SUBMIT ANSWER
Answer #1

Direct materials:

Total materials per unit = Actual quantity of materials + Average waste and spoilage per unit

= 3.20 pounds + 0.10 pound = 3.3 pounds

Material cost before discount = 3.3 pounds*$2 = $6.6

Material cost after discount = $6.6*(1 - 0.05) = $6.6*0.95 = $6.27

Freight cost = 3.3 pounds*0.25 = $0.825

Total direct materials cost per unit = $6.27 + $0.825 = $7.095

Direct labor:

Direct labor hours per unit = 2.10 hours + 0.10 hours = 2.2 hours

Total direct labor cost per unit = 2.2*($13.2 + $2.2) = 2.2*$15.4 = $33.88

Manufacturing overhead:

Total Direct labor hours per unit = 2.2 hours

Manufacturing overhead cost = 2.2 hours*$4.80 = $10.56

Total standard cost per unit = Total direct materials cost per unit + Total direct labor cost per unit + Manufacturing overhead cost

= $7.095 + $33.88 + $10.56 = $51.535 i.e. $51.54